

********************************************************************
UARS WEEKLY STATUS REPORT - 21 JULY 2000
********************************************************************
***********
SOLSTICE
***********
This week, we have seen some degraded solar pointing performance (3-5
arc-minutes vs. 1.5 arc-minutes nominally) but this is still within our
calibration. Additionally, we continue to gather enough data to meet all
the SOLSTICE science objectives during this period of reduced tape
recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans scheduled
for periods of predicted telemetry.
We are preparing for our 4 lamp batch of yearly calibration scans.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
At the present time, the PEM is not active...may UARS get better soon.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
***********
HALOE
***********
HALOE had been operating successfully until the scheduled power off
beginning-of-day July 19,2000. All engineering parameters were within
nominal operating region and the science data processed results are of
excellent quality.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to assist with analysis of the impact
resulting from the failure of UARS Fixed Head Star Tracker 2.
UARS spacecraft attitude control and knowledge are now
being maintained from Fine Sun Sensor data and periodic
operation of Fixed Head Star Tracker 1.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
Continue to notify all PI Software Representatives when Data Gaps
occur in the LEVEL0 data due to the UARS Spacecraft Anomaly.
Data created in the month of May, 2000 will be previewed and loaded
into the Daac Transfer Tables July, 2000.
UARS CDHF REHOST:
- Intentions are on July 31,2000 to begin running current day processing,
Level 0 and NRT Ingest on both the present production and Rehost systems.
Approximate duration 2 weeks. We would then like PI/SW reps to use
the Rehost system the primary system for LO ,NRT and current day processing.
PAC and Production will notify PI software reps regarding data transfer and
account access on the Rehost.
Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, HRDI,
MLS, SUSIM, Solstice, and WINDII.
Built SOLAR cataloging files on the rehost cluster, with MPG user SOL.
Many subsystems on the rehost have recently been re-linked and tested.
A re-linking of all the PI software tested is now required, and in
progress. Correlative suppliers who had already linked to the old images,
have been notified to re-compile and re-link as well.
***********
FOT
***********
REPORTING PERIOD
This report covers 08 July 2000 (Orbit 48227, GMT Day 190) through 14 July 2000
(Orbit 48331 , GMT Day 196).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 192/1311 48265 POWER ON
The beta angle ranged between the angles of 10.6 degrees to 38.5 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+76 degrees on 25 July 2000 (DOY 207).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.3 degrees C. Additional battery performance data for each
battery on 07 July 2000 and 14 June 2000 are provided in Appendix D.
The clock error ranged from -11.7 to 11.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 34
% HALOE = 01
% SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/189 show the observatory data loss to be 296 hours, 37
minutes, 2.841 seconds (an increase of 1 minute 27.488 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
The increase of 1 minute 27.488 seconds occurred over two (2) days:
- 0 minutes 00.576 seconds on Day 189 due a data hit,
- 1 minutes 26.912 seconds on Day 193 due to an unknown reason.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The investigation and correction of the FHST-2 anomaly continued this week. It
appears from the trending data and from a series of tests that FHST-2
experienced an electrical failure which has rendered the tracker inoperable.
The exact nature of the failure is still under investigation. The FOT returned
FHST-1 to operations and powered off FHST-2. FHST-1 is operating at a reduced
capacity but will give UARS a stable attitude until it can be recalibrated.
The FOT has delayed the testing of the new TMONs until the FHST-2 anomaly is
resolved. No time frame has been set as to when testing will resume.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended meetings to discuss FHSTs 1 & 2 on 10 & 12 July 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for August 14, 2000 (DOY 227). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 July 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48230 190/0608 00-316 NBTR NOOPS D190
48235 190/1307 00-319 SUSIM EXCEPTION CLEAR
48236 190/1549 00-320 FHST TESTING
48236 190/1631 00-322 FHST TESTING
48238 190/1739 00-316 NBTR NOOPS D190
48238 190/1739 00-323 FHST TESTING
48239 190/1925 00-321 NBTR NOOPS D191
48239 190/1925 00-324 FHST TESTING
48241 190/2246 00-325 OBC Clk Adj 0.0 to -20.6 msec/day
48245 191/0610 00-321 NBTR NOOPS D191
48255 191/2149 00-327 OBC Clk Adj -20.6 to 0.0 msec/day
48260 192/0544 00-326 NBTR NOOPS D192
48264 192/1228 00-328 OBC Clk Adj 0.0 to 20.6 msec/day
48265 192/1311 00-329 HALOE POWER ON
48268 192/1816 00-326 NBTR NOOPS D192
48268 192/1816 00-330 NBTR NOOPS D193
48275 193/0548 00-330 NBTR NOOPS D193
48280 193/1413 00-332 FHST1 POWER ON
48281 193/1452 00-332 FHST1 POWER OFF
48282 193/1633 00-333 OBC Clk Adj 20.6 to 0.0 msec/day
48283 193/1749 00-330 NBTR NOOPS D193
48283 193/1749 00-331 NBTR NOOPS D194
48290 194/0621 00-331 NBTR NOOPS D194
48298 194/1752 00-331 NBTR NOOPS D194
48298 194/1752 00-334 NBTR NOOPS D195
48305 195/0553 00-334 NBTR NOOPS D195
48309 195/1234 00-335 FHST1 STAR UPDATE
48312 195/1642 00-334 NBTR NOOPS D195
48313 195/1754 00-336 NBTR NOOPS D196
48320 196/0555 00-336 NBTR PLAYBACK
48320 196/0555 00-337 PSU AXIS QB OFF
48321 196/0632 00-336 NBTR NOOPS D196
48323 196/0959 00-338 FHST1 POWER ON
48324 196/1055 00-338 FHST1 POWER OFF
48324 196/1138 00-339 FHST1 POWER ON
48327 196/1643 00-339 FHST1 POWER OFF
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 July 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
ACTEMPANA1 Y-LO TEMP PSU BOARD 48240
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
SUEXCMON1 SUSIM EXCEPTION MONITOR 48232
OBSZ100L Z1 RESIDUAL EXCEEDED 100 ARC/SEC 48325-7
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 July 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
48241 190/22:53 (08 June 00) from 0.00 to -20.6 msec/day
48255 191/22:04 (09 June 00) from -20.6 to 0.00 msec/day
48264 192/12:56 (10 June 00) from 0.00 to 20.6 msec/day
48282 193/16:38 (11 June 00) from 20.6 to 0.00 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 July 2000
07 July 2000 - GMT Day 189 Beta = 10.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------ ------
1 +44.8/+28.0 -3.68/-4.94 +0.0/+0.0 15.5 0.000 0.0
2 +11.2/-67.2 +4.51/+3.87 +28.8/-13.6 26.1 1.040 19.8
3 +28.0/-22.4 +3.55/+2.60 +28.0/-13.2 25.9 1.040 19.3
14 July 2000 - GMT Day 196 Beta = 38.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +56.0/+33.6 -3.05/-4.31 +0.0/+0.0 15.5 0.000 0.0
2 +50.4/-56.0 +5.79/+4.51 +28.0/-15.2 25.0 1.030 22.0
3 +22.4/-33.6 +5.15/+3.23 +27.2/-16.4 25.0 1.030 22.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 July 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
00-031 FOT 48285 193/2111 14170 Generic Late Acq.
00-031 FOT 48317 196/0117 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
00-030 FOT 48211 188/2242 22631 Intermittent drop outs
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-026 FOT see list June 00 14170 Generic Late Acq.
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/28